OpenStudy (anonymous):

Simplify the expression (-1+6i)+(-4+2i) a.5-8i c.-5+8i b.5-2i d.3i

OpenStudy (anonymous):

d. 3i

OpenStudy (anonymous):

That is wrong it should be : C. -5 + 8i

OpenStudy (anonymous):

Tdabboud, -1+6i = 5i, -4+2i = -2i, 5i + -2i = 3i

OpenStudy (unklerhaukus):

(-1+6i)+(-4+2i) = -1 +6i -4+2i = -5 +8i

OpenStudy (anonymous):

you cannot simply add the i like that. i is an imaginary number so it can only be added with other i's. Not like you have done chaplaps

OpenStudy (lilg132):

C

OpenStudy (lilg132):

-1 + -4 = -5 6i + 2i = 8i -5 + 8i

OpenStudy (anonymous):

hmm okay sorry

OpenStudy (anonymous):

How would u rewrite this in vertex form y=x^2+10x+16 a.y=(x+5^2)+41 c.y=(x+10^2)+11 b.y=(x+10^2)-9 d.y=(x+5^2)-9

OpenStudy (anonymous):

The answer is D.

OpenStudy (anonymous):

can u show me how u did it?